Circular No. 9066

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 L18 (SKIFF) G. Sostero, E. Guido, P. Camilleri and E. Prosperi report the recovery

of comet P/2002 S1 (cf. IAUC 7972) from the co-addition of forty unfiltered 60-s CCD exposures obtained remotely on June 15.6 UT with the 0.35-m f/7 reflector at the Skylive-Grove Creek Observatory (near Trunkey, NSW, Australia). The recovery was confirmed from the co-addition of ten unfiltered 300-s exposures with the same equipment on Aug. 18.6. The object was of stellar appearance on each occasion (the first observations on each night are tabulated below); the complete set of astrometry appears on CBET 1913, along with corrected and new pre-discovery observations from Palomar, orbital elements for epoch 2002 Mar. 27, residuals, and an ephemeris.

2009 UT α_{2000} δ_{2000} Mag. Observer
June 15.60811 20^{h}49^{m}13.^{s}30 −49^{o}02′32′′.4 20.0 Sostero et al.
Aug. 18.57587 19 45 30.32 −48 51 00.9 19.6  _{”}

The indicated corrections to the predictions on MPC 59600 and in the 2008/2009 Comet Handbook are ΔT = −0.16 and −0.19 day, respectively. The following improved orbital elements by B. G. Marsden satisfy 149 observations, 2001–2009 (mean residual 0′′.6):

Epoch = 2010 Sept. 1.0 TT

T = 2010 Aug. 14.5215 TT  ω = 37.^{o}8480
e = 0.416859  Ω = 346.8250 }2000.0
q = 2.420072 AU  i = 27.0551

a = 4.150062 AU n^{o} = 0.1165795 P = 8.45 years

THE EDGAR WILSON AWARD 2009 The Smithsonian Astrophysical Observatory announces that the 2009

Edgar Wilson Award for the discovery of comets (cf. IAUCs 6936, 8962) is being divided among the following five individuals for five different comets: Robert E. Holmes, Jr. (Charleston, IL, U.S.A.) for C/2008 N1 (cf. IAUC 8959); Stanislav Matičič (Crni Vrh Observatory, Slovenia) for C/2008 Q1 (cf. IAUC 8966); Michel Ory (Delemont, Switzerland) for P/2008 Q2 (cf. IAUC 8967); Koichi Itagaki (Yamagata, Japan) for C/2009 E1 (cf. IAUC 9026); and Dae-am Yi (Yeongwol-kun, Gangwon-do, Korea) for C/2009 F6 (cf. IAUC 9035).

2009 August 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green